# Providence, USA

### v1

@@ 3.1.1 Future Projections
Figure 8: Similar to Fig. 7 but the projected model-independent dark shower limits from dimuon scouting at the HL-LHC. The different colors correspond to $m_{\hat{\pi}}$ benchmark values, while the light-shaded curves corresponds to current limits in Fig. 7. The uncertainties due to varying trigger efficiency scenarios are shown in bands, with the solid contours in the center stand for limits in the median trigger efficiency scenario in App. A. The 2DV limits are also plotted as dashed curves, corresponding to the median trigger efficiency scenario with uncertainties shown as bands. On the label of the vertical axis, the power of the dimuon BR $n=1(2)$ for 1(2)DV limits.
In the scouting search, the macroscopic decay length of dark mesons is the primary feature that separates the signal from SM backgrounds. As the chance of finding two DVs in an event is suppressed by the single DV efficiency twice, the LHC limit from Run 2 is not as competitive. However, searching for more than one DVs may become possible at a higher luminosity. The excellent dimuon mass resolution referred to earlier implies that, even in the HL-LHC era, events featuring two dimuon DVs that fall into the same mass window will be exceedingly rare, thus legitimizing the background-free assumption. To derive the 2DV constraints on HL-LHC from data scouting conservatively, we first assume the overall efficiency of the event is the product of two individual DVs’ efficiencies. [^5] Very often, in events having a dark shower, there are more than two dark hadrons that decay to the dimuon final state. Only the two DVs with the highest trigger efficiencies are considered. Numerically, we found this approach to be a good approximation since the DV efficiencies in an event are often in a hierarchical structure, dominated by the ones with proper lifetime and high $p_{T}$. In addition, all DVs with $l_{xy}<1$ cm are dropped to avoid introducing flavored backgrounds. We plot the projected HL-LHC 2DV constraints on all benchmarks in Fig. 8 as dashed curves, which need to be scaled by BR$(\hat{\pi}\to\mu^{+}\mu^{-})^{2}$. As expected, the 2DV constraints derived this way are most relevant when the dark hadron’s lifetime makes their trigger efficiencies optimal. The constraints weaken much faster than the 1DV case for both smaller and larger lifetimes. Notably, the difference between 1DV and 2DV $Z^{\prime}$ initiated dark shower limits is much smaller than in the other cases, which is probably caused by the low $N_{\hat{\pi}}$ and more evenly distributed $p_{T}$ among dark pions.
@@ 3.2 Dark Shower Sensitivity at LHCb
Figure 9: Model-independent dark shower limits from recasting current LHCb dimuon DV search LHCb:2020ysn . The format is similar to Fig. 7.
@@ 3.2.1 Future Projections
Owing to an even lower trigger threshold starting from Run 3 LHCbCollaboration:2014vzo and the comparatively low pile-up level achieved, we expect the projected dark shower constraints from dimuon DVs for LHCb in run 3 or the high luminosity phase to become stronger than a mere luminosity rescaling. To project the dark shower limit in the high luminosity era, we follow the approach described in Ref. Ilten:2016tkc . In particular, the cut on the dimuon DV is softened so that both muon tracks only need to have $p_{T}>0.5$ GeV, $|p|>10$ GeV, and $\eta\in[2,5]$. The DVs also need to satisfy its total $p_{T}>1$ GeV and $l_{xy}\in[6,22]$ mm. The future LHCb background in a $m_{\mu\mu}$ window is then rescaled from that of Ref. LHCb:2020ysn with the same magnification factor as the signal efficiency improvement. In addition, due to a potential $m_{\mu\mu}$ resolution improvement, the $m_{\mu\mu}$ mass window width is taken to be 16 MeV ($0.016\times m_{\mu\mu}$) when $m_{\mu\mu}\leqslant(>)$1 GeV Ilten:2016tkc , which further decreases the background. The background deduced this way is compatible with Ref. Ilten:2016tkc . Finally, a global detector efficiency factor of 0.7 is applied universally for all benchmark signals. The resulting 1DV projections are shown in Fig. 10.
@@
We commence with the exclusive decays to $K^{+}K^{-}$, applicable to Higgs portal decays Cheng:2021kjg ; CidVidal:2019urm or vector dark meson decays Ilten:2018crw . Dark shower signal samples are simulated using the same method as before, assuming $\hat{\pi}\to K^{+}K^{-}$ decays via the Higgs portal. We select benchmark dark pion masses to avoid overlap with the abundant $\phi$, $\Lambda^{0}$, and $D^{0}$ backgrounds. The selection rules in Ref. CidVidal:2019urm are applied at the truth level. Specifically, both $K^{\pm}$ tracks must satisfy $52<\eta<$ 5, $p_{T}>0.5$ GeV, and an impact parameter from the PV greater than $0.1$ mm. Furthermore, each reconstructed $KK$ DV must have $l_{xy}\in[6,25]$ mm, the longitudinal displacement $l_{z}<400$ mm, $p_{T}>10$ GeV, and an impact parameter from the PV less than $0.1$ mm. Two signal regions are defined by the DV’s $l_{xy}$, namely $l_{xy}\in[6,10]$ mm and $l_{xy}\in[14,25]$ mm CidVidal:2019urm . Similar to previous analyses, no isolation criteria are imposed. A global detector efficiency factor of 0.7 is considered to account for detector effects. We obtain the combinatorial background level from Ref. CidVidal:2019urm , which employs inclusive Pythia simulation. For non-isolated $KK$ DVs, the expected background in the range $m_{KK}\in[1,2]$ GeV is approximately $\mathcal{O}(10^{5}-10^{6})$ when $L=15$ ab-1, depending on the $l_{xy}$ bin. Such a high combinatorial level aligns with LHCb $D^{0}\to K^{+}K^{-}$ studies LHCb:2019dom ; LHCb:2022vcc . The substantial backgrounds diminish the search’s sensitivity to signal events and render the result susceptible to systematic effects. By requiring two DVs with identical mass in a single event, the number of combinatorial backgrounds dramatically reduces by many orders of magnitude at the cost of lower signal yields CidVidal:2019urm . To define the 2DV signal region, two $K^{+}K^{-}$ DV satisfying all kinematic requirements above are required to have $l_{xy}\in[14,25]$ mm. [^7] The requirement corresponds to the signal region $d_{1}$ in Ref. CidVidal:2019urm . The result 2DV limits at the HL-LHC era are plotted in Fig. 11.
Figure 11: The projected LHCb dark shower limits with $\hat{\pi}\to K^{+}K^{-}$ decays at the HL-LHC era. Here only $m_{\hat{\pi}}=1.2$ and 2 GeV benchmarks above 1 GeV are shown, while the background size largely drives the differences between 1DV limits. The 2DV limits from varying $m_{\hat{\pi}}$, on the contrary, are much closer.
@@ 3.3 Dark Shower Sensitivity at Proposed Auxiliary Detectors
Dark hadrons with proper decay lengths $\gtrsim\mathcal{O}$(m) will predominantly decay outside the multipurpose detector systems previously mentioned, resulting in reduced sensitivities which are apparent in, $e.g.$, Fig. 7. In such cases, LHC auxiliary detectors targeting LLPs can play a crucial role in dark shower searches due to their sensitivity to longer lifetime decays Aielli:2019ivi ; Curtin:2018mvb ; Liebersbach:2024kzc . Here we present model-independent limits for dark showers, which may be used generally.
@@ 3.4 Future High-Intensity Experiments
The $e^{+}e^{-}\to Z\to$ dark shower signal samples are generated at the $Z$ pole ($\sqrt{s}=m_{Z}$). We make the following assumptions in our analysis. The dimuon DV signal is recognized if both $\mu$ satisfy $p_{T,\mu}>0.5$ GeV, $|p_{\mu}|>10$ GeV and $|\eta_{\mu}|<5$. In addition, the muon pair forms a DV and needs to have $p_{T,\mu\mu}>2$ GeV and $|\eta_{\mu}|<5$. A DV must have its invariant mass within the window of $[0.99~{}m_{\hat{\pi}},1.01m_{\hat{\pi}}]$ and its transverse displacement ł$\l_{xy}\in[0.5,100]$ cm. Finally, the signal yield is obtained by applying a global detector efficiency of 0.7 on each vertex. We also assume that the background is negligible for such well-reconstructed dimuon DV. Fig 13 shows the projected 95% C.L. exclusion limit on BR($Z\to\psi\bar{\psi}$)$\times$×BR$(\hat{\pi}\to\mu^{+}\mu^{-})$ at the Tera-$Z$, corresponding to $10^{12}$ on-shell $Z$ decays. The optimal limit is achieved when $c\tau(\hat{\pi})\in[0.1-10]$ cm. Within this optimal range, the signal efficiency is high enough, together with a very low background level expected, leading to a best reach of BR($Z\to\psi\bar{\psi}$)$\times$×BR($\hat{\pi}\to\mu^{+}\mu^{-}$) close to the inverse of the total $Z$ number ($10^{-12}$).
@@
4 Model-Independent Limits on Dark Hadrons Produced by FCNC Meson Decays
@@ 4.1 FCNC Constraints and Projections at CMS and LHCb
Since the FCNC decay rates to dark hadrons are suppressed by $f_{a}$, off-diagonal CKM matrix elements, and loop factors, we expect no more than one dark pion to be produced per FCNC decay. [^9] Since the dark QCD could also shower after the dark fermion pair is produced, it is also possible that more than one dark hadron are created by the FCNC decay, $e.g.$, $B\to K+n\hat{\pi}$ process. In this work, we only consider the case of single dark pion production. The signal feature at high-energy colliders will be one DV produced along the direction of a heavy-flavor initiated jet. In this case, the accompanying SM particle like $K^{0}$ or $\pi^{0}$ from FCNC decay may not be reconstructed. Meanwhile particles from the rest of the jet or meson decays make the dimuon DV unisolated.101010 When the dark hadron is light enough that it is still relativistic in the $B$ meson’s rest frame, we expect the kinematics from $B\to X_{s}\hat{\pi}$ to be similar to the exclusive $B\to K^{(\ast)}\hat{\pi}$ decays. Consequently, our model-independent strategy is potentially applicable to inclusive FCNC decays.
Figure 14: Current and projected (HL-)LHC limits on dark hadrons from FCNC $b$-hadron decays based on the CMS dimuon DV scouting strategy. The solid curves and the corresponding band stands for the HL-LHC projections and uncertainties, while the current limit is shown as dashed curves.
@@
Figure 15: $B$ meson FCNC decay to dark hadron limits given by LHCb. Dashed curves are currents
@@ 4.2 FCNC Constraints at Auxiliary Detectors at HL-LHC
Figure 16: Model-independent limits on exclusive BR($B\to\hat{\pi}K$) at various auxiliary detectors. The two sets of MATHUSLA curves correspond to the original and updated designs respectively.
@@ 4.3 FCNC Constraints from Other Experiments
For bounds of this type, we focus on the CHARM experiment CHARM:1985anb . We follow the recast in Ref. NA62:2023nhs using muon pair final states. [^12] We thank Reuven Balkin for pointing out the proper reference here. For different dark hadron mass benchmarks, the model-independent results are shown in Fig. 17, with the maximum sensitivity reached for dark hadron lifetimes of $\sim\mathcal{O}(10)$ m.
Belle II: Lepton collider experiments operating at $\sqrt{s}\simeq 10$ GeV are also known as $B$ factories as $\Upsilon$(4S) can be produced on resonance, which will promptly decay to two $B$ mesons with large cross section. The leading $B$ factory for the next decades is Belle II Kou:2018nap . For the projection at Belle II, we adopt the method in Ref. Filimonova:2019tuy . The $B$ mesons from $\Upsilon(4S)$ decays have a fixed boost of $\beta\gamma\simeq 0.284$ along the beam axis, while their transverse boosts are negligible. To reconstruct the DV, both tracks need to leave enough hits to be reconstructed and form DVs Bernreuther:2022jlj . The signal efficiency and the effective volume of the Belle II detector become smaller than the detector’s geometric size, with negligible backgrounds Ferber:2022ewf . Here, we adopt a simplified detector geometry of the Belle II vertex detector and central drift chamber Kou:2018nap ; Ferber:2022ewf . All dark pions that decay inside or before reaching the vertex detector or the central drift chamber with $0.260>l_{xy}>$ 0.2 cm are kept. The macroscopic $l_{xy}$ and a narrow resonance peak can reduce combinatorial backgrounds to a negligible level for final states with only two charged tracks, i.e., $\mu^{+}\mu^{-}$, $\pi^{+}\pi^{-}$, or $K^{+}K^{-}$. [^13] For electron pair final states, further cuts are needed to remove large backgrounds from converted photons. Assuming a signal efficiency of $\sim 1$ and negligible backgrounds, the 95% C.L. exclusion limit then overlaps with the $N_{\rm sig}=3$ contour in Fig. 17. The experiment could probe an exclusive FCNC $B\to K+\hat{\pi}$ decay with a branching ratio down to $\mathcal{O}(10^{-10})$ level when $c\tau(\hat{\pi})$ is between 0.1-10 cm.

### v2

@@ 3.1.1 Future Projections
Figure 8: Similar to Fig. 7 but the projected dark shower limits from dimuon scouting at the HL-LHC. The different colors correspond to $m_{\hat{\pi}}$ benchmark values, while the light-shaded curves corresponds to current limits in Fig. 7. The uncertainties due to varying trigger efficiency scenarios are shown in bands, with the solid contours in the center stand for limits in the median trigger efficiency scenario in App. A. The 2DV limits are also plotted as dashed curves, corresponding to the median trigger efficiency scenario with uncertainties shown as bands. On the label of the vertical axis, the power of the dimuon BR $n=1(2)$ for 1(2)DV limits.
In the scouting search, the macroscopic decay length of dark mesons is the primary feature that separates the signal from SM backgrounds. As the chance of finding two DVs in an event is suppressed by the single DV efficiency twice, the LHC limit from Run 2 is not as competitive. However, searching for more than one DVs may become possible at a higher luminosity. The excellent dimuon mass resolution referred to earlier implies that, even in the HL-LHC era, events featuring two dimuon DVs that fall into the same mass window will be exceedingly rare, thus legitimizing the background-free assumption. To derive the 2DV constraints on HL-LHC from data scouting conservatively, we first assume the overall efficiency of the event is the product of two individual DVs’ efficiencies. [^7] Very often, in events having a dark shower, there are more than two dark hadrons that decay to the dimuon final state. Only the two DVs with the highest trigger efficiencies are considered. Numerically, we found this approach to be a good approximation since the DV efficiencies in an event are often in a hierarchical structure, dominated by the ones with proper lifetime and high $p_{T}$. In addition, all DVs with $l_{xy}<1$ cm are dropped to avoid introducing flavored backgrounds. Based on the assumption of background and Poisson statistics, the 95% C.L. limit of a background-free signal region corresponds to a signal yield $\sim$ 3. We plot the projected 95% C.L. HL-LHC 2DV constraints on all benchmarks in Fig. 8 as dashed curves, which need to be scaled by BR$(\hat{\pi}\to\mu^{+}\mu^{-})^{2}$. As expected, the 2DV constraints derived this way are most relevant when the dark hadron’s lifetime makes their trigger efficiencies optimal. The constraints weaken much faster than the 1DV case for both smaller and larger lifetimes. Notably, the difference between 1DV and 2DV $Z^{\prime}$ initiated dark shower limits is much smaller than in the other cases, which is probably caused by the low $N_{\hat{\pi}}$ and more evenly distributed $p_{T}$ among dark pions.
@@ 3.2 Dark Shower Sensitivity at LHCb
Figure 9: The 95% C.L. dark shower limits from recasting current LHCb dimuon DV search LHCb:2020ysn . The format is similar to Fig. 7.
@@ 3.2.1 Future Projections
Owing to an even lower trigger threshold starting from Run 3 LHCbCollaboration:2014vzo and the comparatively low pile-up level achieved, we expect the projected dark shower constraints from dimuon DVs for LHCb in run 3 or the high luminosity phase to become stronger than a mere luminosity rescaling. To project the dark shower limit in the high luminosity era, we follow the approach described in Ref. Ilten:2016tkc . In particular, the cut on the dimuon DV is softened so that both muon tracks only need to have $p_{T}>0.5$ GeV, $|p|>10$ GeV, and $\eta\in[2,5]$. The DVs also need to satisfy its total $p_{T}>1$ GeV and $l_{xy}\in[6,22]$ mm. The future LHCb background in a $m_{\mu\mu}$ window is then rescaled from that of Ref. LHCb:2020ysn with the same magnification factor as the signal efficiency improvement. In addition, due to a potential $m_{\mu\mu}$ resolution improvement, the $m_{\mu\mu}$ mass window width is taken to be 16 MeV ($0.016\times m_{\mu\mu}$) when $m_{\mu\mu}\leqslant(>)$1 GeV Ilten:2016tkc , which further decreases the background. The background deduced this way is compatible with Ref. Ilten:2016tkc . Finally for the detector efficiency, in a previous study (Ref. Cheng:2021kjg ) it was found that a detector efficiency in the range $[0.4,0.8]$ can reproduce the LHCb results LHCb:2020ysn . For simplicity, we apply a global detector efficiency factor of 0.7 for all benchmark signals, with the optimistic hope that it will not be worse at the future LHCb runs. It is also comparable to though slightly higher than the one used in Ref. Ilten:2016tkc . The resulting 1DV projections are shown in Fig. 10.
@@
We commence with the exclusive decays to $K^{+}K^{-}$, applicable to Higgs portal decays Cheng:2021kjg ; CidVidal:2019urm or vector dark meson decays Ilten:2018crw . Dark shower signal samples are simulated using the same method as before, assuming $\hat{\pi}\to K^{+}K^{-}$ decays via the Higgs portal. We select benchmark dark pion masses to avoid overlap with the abundant $\phi$, $\Lambda^{0}$, and $D^{0}$ backgrounds. The selection rules in Ref. CidVidal:2019urm are applied at the truth level. Specifically, both $K^{\pm}$ tracks must satisfy $52<\eta<$ 5, $p_{T}>0.5$ GeV, and an impact parameter from the PV greater than $0.1$ mm. Furthermore, each reconstructed $KK$ DV must have $l_{xy}\in[6,25]$ mm, the longitudinal displacement $l_{z}<400$ mm, $p_{T}>10$ GeV, and an impact parameter from the PV less than $0.1$ mm. Two signal regions are defined by the DV’s $l_{xy}$, namely $l_{xy}\in[6,10]$ mm and $l_{xy}\in[14,25]$ mm CidVidal:2019urm . Similar to previous analyses, no isolation criteria are imposed. A global detector efficiency factor of 0.7 is considered to account for detector effects. [^9] The choice is more conservative than the efficiency adopted in Ref. CidVidal:2019urm . We obtain the combinatorial background level from Ref. CidVidal:2019urm , which employs inclusive Pythia simulation. For non-isolated $KK$ DVs, the expected background in the range $m_{KK}\in[1,2]$ GeV is approximately $\mathcal{O}(10^{5}-10^{6})$ when $L=15$ ab-1, depending on the $l_{xy}$ bin. Such a high combinatorial level aligns with LHCb $D^{0}\to K^{+}K^{-}$ studies LHCb:2019dom ; LHCb:2022vcc . The substantial backgrounds diminish the search’s sensitivity to signal events and render the result susceptible to systematic effects. By requiring two DVs with identical mass in a single event, the number of combinatorial backgrounds dramatically reduces by many orders of magnitude at the cost of lower signal yields CidVidal:2019urm . To define the 2DV signal region, two $K^{+}K^{-}$ DV satisfying all kinematic requirements above are required to have $l_{xy}\in[14,25]$ mm. [^10] The requirement corresponds to the signal region $d_{1}$ in Ref. CidVidal:2019urm . The result 2DV limits at the HL-LHC era are plotted in Fig. 11.
Figure 11: The projected LHCb 95% C.L. dark shower limits with $\hat{\pi}\to K^{+}K^{-}$ decays at the HL-LHC era. Here only $m_{\hat{\pi}}=1.2$ and 2 GeV benchmarks above 1 GeV are shown, while the background size largely drives the differences between 1DV limits. The 2DV limits from varying $m_{\hat{\pi}}$, on the contrary, are much closer.
@@ 3.3 Dark Shower Sensitivity at Proposed Auxiliary Detectors
Dark hadrons with proper decay lengths $\gtrsim\mathcal{O}$(m) will predominantly decay outside the multipurpose detector systems previously mentioned, resulting in reduced sensitivities which are apparent in, $e.g.$, Fig. 7. In such cases, LHC auxiliary detectors targeting LLPs can play a crucial role in dark shower searches due to their sensitivity to longer lifetime decays Aielli:2019ivi ; Curtin:2018mvb ; Liebersbach:2024kzc . Here we present limits for dark showers, which may be used generally.
@@ 3.4 Future High-Intensity Experiments
The $e^{+}e^{-}\to Z\to$ dark shower signal samples are generated at the $Z$ pole ($\sqrt{s}=m_{Z}$). We make the following assumptions in our analysis. The dimuon DV signal is recognized if both $\mu$ satisfy $p_{T,\mu}>0.5$ GeV, $|p_{\mu}|>10$ GeV and $|\eta_{\mu}|<5$. In addition, the muon pair forms a DV and needs to have $p_{T,\mu\mu}>2$ GeV and $|\eta_{\mu}|<5$. A DV must have its invariant mass within the window of $[0.99~{}m_{\hat{\pi}},1.01m_{\hat{\pi}}]$ and its transverse displacement ł$\l_{xy}\in[0.5,100]$ cm. Finally, the signal yield is obtained by applying a global detector efficiency of 0.7 on each vertex as the benchmark value. Here we take the conservative assumption that the detector efficiency at future $Z$-factories will be no worse than its LHCb counterpart. We also assume that the background is negligible for such well-reconstructed dimuon DV. Fig 13 shows the projected 95% C.L. exclusion limit on BR($Z\to\psi\bar{\psi}$)$\times$×BR$(\hat{\pi}\to\mu^{+}\mu^{-})$ at the Tera-$Z$, corresponding to $10^{12}$ on-shell $Z$ decays. The optimal limit is achieved when $c\tau(\hat{\pi})\in[0.1-10]$ cm. Within this optimal range, the signal efficiency is high enough, together with a very low background level expected, leading to a best reach of BR($Z\to\psi\bar{\psi}$)$\times$×BR($\hat{\pi}\to\mu^{+}\mu^{-}$) close to the inverse of the total $Z$ number ($10^{-12}$).
@@
4 Limits on Dark Hadrons Produced by FCNC Meson Decays
@@ 4.1 FCNC Constraints and Projections at CMS and LHCb
Since the FCNC decay rates to dark hadrons are suppressed by $f_{a}$, off-diagonal CKM matrix elements, and loop factors, we expect no more than one dark pion to be produced per FCNC decay. [^12] Since the dark QCD could also shower after the dark fermion pair is produced, it is also possible that more than one dark hadron are created by the FCNC decay, $e.g.$, $B\to K+n\hat{\pi}$ process. In this work, we only consider the case of single dark pion production. The signal feature at high-energy colliders will be one DV produced along the direction of a heavy-flavor initiated jet. In this case, the accompanying SM particle like $K^{0}$ or $\pi^{0}$ from FCNC decay may not be reconstructed. Meanwhile particles from the rest of the jet or meson decays make the dimuon DV unisolated.131313 When the dark hadron is light enough that it is still relativistic in the $B$ meson’s rest frame, we expect the kinematics from $B\to X_{s}\hat{\pi}$ to be similar to the exclusive $B\to K^{(\ast)}\hat{\pi}$ decays. Consequently, our strategy is potentially applicable to inclusive FCNC decays.
Figure 14: Current and projected (HL-)LHC 95% C.L. limits on dark hadrons from FCNC $b$-hadron decays based on the CMS dimuon DV scouting strategy. The solid curves and the corresponding band stands for the HL-LHC projections and uncertainties, while the current limit is shown as dashed curves.
@@
Figure 15: The 95% C.L. $B$ meson FCNC decay to dark hadron limits given by LHCb. Dashed curves are current ones.
@@ 4.2 FCNC Constraints at Auxiliary Detectors at HL-LHC
Figure 16: The 95% C.L. limits on exclusive BR($B\to\hat{\pi}K$) at various auxiliary detectors. The two sets of MATHUSLA curves correspond to the original and updated designs respectively.
@@ 4.3 FCNC Constraints from Other Experiments
For bounds of this type, we focus on the CHARM experiment CHARM:1985anb . We follow the recast in Ref. NA62:2023nhs using muon pair final states. [^15] We thank Reuven Balkin for pointing out the proper reference here. For different dark hadron mass benchmarks, the results are shown in Fig. 17, with the maximum sensitivity reached for dark hadron lifetimes of $\sim\mathcal{O}(10)$ m.
Belle II: Lepton collider experiments operating at $\sqrt{s}\simeq 10$ GeV are also known as $B$ factories as $\Upsilon$(4S) can be produced on resonance, which will promptly decay to two $B$ mesons with large cross section. [^16] One may also consider dark shower signals at $B$ factories. However, in our EW portal benchmarks, the mediating bosons are too heavy to be produced on-shell at $B$ factories and leading to suppressed signal yields. We found no particular advantage of dark shower search at $B$ factories compared to the LHC. For dark shower searches at Belle II induced by other portal interactions, see Ref. Bernreuther:2022jlj . The leading $B$ factory for the next decades is Belle II Kou:2018nap . For the projection at Belle II, we adopt the method in Ref. Filimonova:2019tuy . The $B$ mesons from $\Upsilon(4S)$ decays have a fixed boost of $\beta\gamma\simeq 0.284$ along the beam axis, while their transverse boosts are negligible. To reconstruct the DV, both tracks need to leave enough hits to be reconstructed and form DVs Bernreuther:2022jlj . The signal efficiency and the effective volume of the Belle II detector become smaller than the detector’s geometric size, with negligible backgrounds Ferber:2022ewf . Here, we adopt a simplified detector geometry of the Belle II vertex detector and central drift chamber Kou:2018nap ; Ferber:2022ewf . All dark pions that decay inside or before reaching the vertex detector or the central drift chamber with $0.260>l_{xy}>$ 0.2 cm are kept. The macroscopic $l_{xy}$ and a narrow resonance peak can reduce combinatorial backgrounds to a negligible level for final states with only two charged tracks, i.e., $\mu^{+}\mu^{-}$, $\pi^{+}\pi^{-}$, or $K^{+}K^{-}$. [^17] For electron pair final states, further cuts are needed to remove large backgrounds from converted photons. Assuming a signal efficiency of $\sim 1$ and negligible backgrounds, the 95% C.L. exclusion limit then overlaps with the $N_{\rm sig}=3$ contour in Fig. 17. The experiment could probe an exclusive FCNC $B\to K+\hat{\pi}$ decay with a branching ratio down to $\mathcal{O}(10^{-10})$ level when $c\tau(\hat{\pi})$ is between 0.1-10 cm.